OpenStudy (naruko885):

A cylinder and a cone have the same diameter: 8 inches. The height of the cylinder is 5 inches. The height of the cone is 15 inches. Use π = 3.14. What is the relationship between the volume of this cylinder and this cone? Explain your answer by determining the volume of each and comparing them. Show all your work.

OpenStudy (naruko885):

I could really use some help on this..

OpenStudy (wanell):

im gonna run through it with you

OpenStudy (naruko885):

ok thank you so much!!

OpenStudy (wanell):

Volume of a cylinder: V=πr2h V=(3.14)(4)2(5) What's 4 * 4?

OpenStudy (naruko885):

16 right?

OpenStudy (wanell):

Yes, and what's 3.14 * 16 * 5?

OpenStudy (naruko885):

251.2 I think.

OpenStudy (wanell):

yes

OpenStudy (wanell):

Volume of a cone: V=13πr2h V=13(3.14)(4)2(15) What's 1/3 * 3.14 * 16 * 15?

OpenStudy (naruko885):

.....oh I get it! thank you so much!

OpenStudy (wanell):

so u got the rest

OpenStudy (naruko885):

I think so but could you tell me if im right.

OpenStudy (wanell):

yea

OpenStudy (naruko885):

oh I think I messed it up could you tell me what I should do now?

OpenStudy (wanell):

the volume of a cone is ONE THIRD the volume of a cylinder.

OpenStudy (wanell):

relationship is that they're the same

OpenStudy (naruko885):

ok so I think i get it. so for the question I just put down the work and then explain the different of a cone and cylinder.

OpenStudy (wanell):

explain the relationship between them

OpenStudy (wanell):

Volume of a cone is 1/3 the volume of a cylinder with the same base. Cone = V= 1/3 Bh where the B = the base area and h is the height Area of a circle = πr² (if the diameter is 8 the radius of this circle is 4) A = π4² = π16 usually notated as 16π (numerical coefficients come first) V= 1/3 x 16 x 18 V= 6 x 16 (I took 1/3 of 18 which is 6) V of the cone = 96 π in³ Volume of a cylinder is = Bh or in this case 16π x 18 = 288π Comparing the two you can see that the cylinder is 3 x the volume of the cone or the cone is 1/3 the volume of the cylinder.

OpenStudy (naruko885):

ok I did it thanks. do you think you could help me with 3 more in a different box?

OpenStudy (wanell):

sure if i can